A sample of NH3 weighs 3.956 g. How many molecules of NH3 are in the sample?


Your answer should be in scientific notation, to three significant figures. Type the letter "x" or "e" in the answer field when you are ready to enter the power of 10.


     molecules of NH3 


1
Expert's answer
2021-04-05T07:42:13-0400

N=n÷NA

Where; N= no of Molecules

n= No of mole

NA= 6.022×10²³

n= Mass/molar mass

Molar mass of NH3 = 14+3 = 17

n = 3.956÷17= 0.2327 mol

N= 0.2327×6.022×10²³

No of Molecules = 1.4×10²³